# Quantum Mycielskians: symmetries, twin vertices and distinguishing labelings

### v1

@@ Abstract.
We study the properties of a generalization of the Mycielski construction to quantum graphs. We discuss quantum symmetries of quantum Mycielskians and their relation to quantum symmetries of the original quantum graphs. We introduce a quantum generalization of the notion of graphs with twin vertices and check that this concept coincides with its classical version for classical graphs with few vertices. Finally, we propose the quantum generalization of the distinguishing number and study its behavior under the quantum Mycielski construction.
@@ 1. Introduction
Classical graph theory has a variety of intriguing applications beyond pure mathematics. Its use in communication theory leads to concepts that can subsequently be generalized into the framework of quantum information theory (see, e.g. [11]). This motivated the notion of quantum graphs defined in terms of Kraus operators associated with quantum communication channels. Several aspects of quantum graphs and their characteristics have recently been widely studied; see, e.g. [6, 7, 11, 13]. Some of the concepts known from the classical graph theory that were studied in the quantum framework are chromatic numbers and clique numbers. The (classical) chromatic number is defined as the minimal number of labels (colors) that can be associated with vertices of a given graph so that the endpoints of any edge do not share the same labeling. The clique number is the maximal size of the complete graph that can be embedded as a subgraph in the original graph. For example, graphs whose clique number is equal to $2$ are triangle-free.
One of the problems in classical graph theory was to determine if it is possible to construct triangle-free graphs with an arbitrarily large chromatic number. In [16] Mycielski has shown that out of a given triangle-free graph $G$, one can construct the new one, denoted by $\mu(G)$, which is also triangle-free, contains the original one as a subgraph, and whose chromatic number is larger by one than the chromatic number of the original graph.
The natural question of extension of this result into the quantum framework has been addressed in [5], where two of the present authors proposed a construction for quantum graphs that mimics the classical construction of Mycielski. Starting from a (finite) quantum graph one can construct a new one, dubbed the quantum Mycielskian, in which the original one can be embedded. The behavior of graph characteristics under this transformation was studied in detail. While the clique number is an invariant of this operation, the (quantum) chromatic numbers satisfy only a weaker property: they can be enlarged by at most one.
In this paper, we study the further properties of this construction in the quantum framework. We begin with Section 1.1 summarizing the classical Mycielski construction. We continue in Section 1.2 with the notation for quantum graphs and the actions of quantum groups on them that will be of great use in the remaining parts of the paper. We will be especially interested in actions that preserve partitions of unity, which in the non-commutative framework correspond to colorings or labelings (see Section 4). This concept is briefly discussed in Section 1.3. At the end of Section 1 we recall the construction of quantum Mycielskian. In Section 2 we discuss isomorphisms of quantum Mycielskians, while Section 3 is devoted to the study of how the Mycielski construction for quantum graphs affects their symmetries. This leads us to the concept of quantum twin vertices that we then study in Sections 3.1 and 3.2 (as well as in the Appendix). Yet another characteristic of classical graphs that can also be defined in the quantum setting is that of distinguishing labeling. We define the distinguishing number in Section  4 and study the impact of the quantum Mycielski construction.
@@ 1.1. Classical Mycielski construction
Let $G=(V_{G},E_{G})$ be a (finite, undirected) graph. The Mycielskian of $G$, denoted by $\mu(G)$ is a graph with the set of vertices $V_{\mu(G)}=\{\bullet\}\sqcup V_{G}\sqcup V_{G}$, where $\bullet$ denotes a distinguished point, called master vertex. The edge set $E_{\mu(G)}$ can be specified by the adjacency matrix $A_{\mu(G)}$. We recall that the adjacency matrix $A_{G}$ for an undirected graph $G$ is a $\{0,1\}$-matrix whose columns and rows are indexed by the set of vertices, and $(A_{G})_{i,j}$ is non-zero if and only if $\{v_{i},v_{j}\}\in E$. For the Mycielskian, we have
$A_{\mu(G)}=\begin{bmatrix}0&\vec{0}^{T}&\vec{1}^{T}\\ \vec{0}&A_{G}&A_{G}\\ \vec{1}&A_{G}&0\end{bmatrix},$
where $\vec{1}$ (resp. $\vec{0}$) is the column vector with all entries being $1$ (resp. 00). Mycielski’s theorem [16] says that denoting by $\omega(\cdot)$ the clique number of a graph we have $\omega(\mu(G))=\omega(G)$, i.e. the clique number remains unchanged under the passage from $G$ to $\mu(G)$, and for the chromatic numbers $\chi$ we have $\chi(\mu(G))=\chi(G)+1$.
@@
$V_{\mu_{r-1}(G)}=\{\bullet\}\sqcup V^{1}\sqcup\ldots\sqcup V^{n}$
@@
$E=E_{G}\cup\bigcup_{i=0}^{r-1}\bigl{\{}\{v^{i}_{j},v^{i+1}_{j^{\prime}}\}\,\bigr{|}\bigl{.}\,\{v^{0}_{j},v^{0}_{j^{\prime}}\}\in E_{G}\bigr{\}}\cup\bigl{\{}\{v^{r}_{1},\bullet\},\{v^{r}_{2},\bullet\},\ldots,\{v^{r}_{|V_{G}|},\bullet\}\bigr{\}}.$
@@ 1.2. Quantum graphs and their quantum symmetries
In what follows, the $\mathrm{C}^{*}$-algebra $\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}})$ will be denoted by $\operatorname{C}(\mathcal{G})$ and the Hilbert space $L^{2}(\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}}),\psi_{\scriptscriptstyle\mathcal{G}})$ by $L^{2}(\mathcal{G})$. We will also shorten $\delta(\psi_{\scriptscriptstyle\mathcal{G}})$ to $\delta_{\scriptscriptstyle\mathcal{G}}$.
Definition 1.2 ([7, Definition 6.3]).
@@
The quantum symmetry group of a quantum graph $\mathcal{G}$ is by construction a quantum subgroup of a well-studied quantum group $\operatorname{Aut}^{+}(\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}}),\psi_{\scriptscriptstyle\mathcal{G}})$ of symmetries of the quantum space $\mathbb{V}_{\scriptscriptstyle\mathcal{G}}$ together with the $\delta$-form $\psi_{\scriptscriptstyle\mathcal{G}}$ (this larger quantum group does not preserve the adjacency matrix). The quantum groups of symmetries of finite quantum spaces equipped with a $\delta$-form were introduced in [2].
@@ 1.3. Actions preserving a partition of unity
Let $\mathsf{A}$ be a unital $\mathrm{C}^{*}$-algebra and let $\rho\colon\mathsf{A}\to\mathsf{A}\otimes\operatorname{C}(\mathbb{G})$ be an action of a compact quantum group $\mathbb{G}$ on the quantum space underlying $\mathsf{A}$. Let $\mathfrak{P}=\{P_{1},\dotsc,P_{n}\}$ be a partition of unity in $\mathsf{A}$. Let $\mathsf{J}$ be the ideal in $\operatorname{C}(\mathbb{G})$ generated by the elements
@@
[…]
@@ 1.4. Mycielskians of quantum graphs
Here we recall the construction introduced in [5]. For $r\geq{2}$ the Mycielski construction of order $r-1$ for quantum graphs is the procedure of passing from a quantum graph $\mathcal{G}$ to a new quantum graph $\mu_{r-1}(\mathcal{G})$ defined as follows ([5, Section 4]):
@@
    $A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\left[\begin{smallmatrix}\lambda\\ x_{1}\\[-2.0pt] \vdots\\[2.0pt] x_{r}\end{smallmatrix}\right]=\left[\begin{smallmatrix}\delta_{\scriptscriptstyle\mathcal{G}}^{2}\psi_{\scriptscriptstyle\mathcal{G}}(x^{r})\\ A_{\scriptscriptstyle\mathcal{G}}(x_{1}+x_{2})\\ A_{\scriptscriptstyle\mathcal{G}}(x_{2}+x_{3})\\[-2.0pt] \vdots\\[2.0pt] A_{\scriptscriptstyle\mathcal{G}}(x_{r-2}+x_{r-1})\\ \lambda\mathds{1}_{\operatorname{C}(\mathcal{G})}+A_{\scriptscriptstyle\mathcal{G}}x_{r-1}\end{smallmatrix}\right].$
@@
for $k=1,\dotsc,r$. Note that although $\iota_{0},\dotsc,\iota_{r}$, and their adjoints are not multiplicative, they are all *-maps. In this notation, we have (cf. [5, Eq. (9)]):
$A_{\mu_{r-1}(\mathcal{G})}=\delta_{\mathcal{G}}\iota_{r}\eta_{\mathcal{G}}\iota_{0}^{\ast}+\delta_{\mathcal{G}}\iota_{0}\eta_{\mathcal{G}}^{\ast}\iota_{r}^{\ast}+\iota_{1}A_{\mathcal{G}}\iota_{1}^{\ast}+\sum\limits_{k=1}^{r-1}\bigl{(}\iota_{k}A_{\mathcal{G}}\iota_{k+1}^{\ast}+\iota_{k+1}A_{\mathcal{G}}\iota_{k}^{\ast}\bigr{)}.$
@@ 2. Quantum isomorphisms of Mycielskians
Let $\mathcal{G}$ and $\mathcal{F}$ be quantum graphs. We say that $\mathcal{G}$ and $\mathcal{F}$ are quantum isomorphic if the quantum groups $\operatorname{QAut}(\mathcal{G})$ and $\operatorname{QAut}(\mathcal{F})$ are monoidally equivalent ([6, Section 4] and [7, Definition 6.5]). The notion of monoidal equivalence of compact quantum groups is described e.g. in [10].
@@
Let $\mathcal{F}$ and $\mathcal{G}$ be quantum graphs. If $\mathcal{F}$ and $\mathcal{G}$ are quantum isomorphic, then $\delta_{\mathcal{F}}=\delta_{\scriptscriptstyle\mathcal{G}}$.
@@
The monoidal equivalence of the quantum groups $\operatorname{QAut}(\mathcal{G})$ and $\operatorname{QAut}(\mathcal{F})$ amounts to a certain algebra $\mathcal{O}(\operatorname{QAut}(\mathcal{G}),\operatorname{QAut}(\mathcal{F}))$ being non-zero ([7, Theorem 6.6]). Since this algebra is a homomorphic image for the corresponding linking algebra for the quantum automorphism groups $\operatorname{Aut}^{+}(\operatorname{C}(\mathbb{V}_{\mathcal{F}}),\psi_{\mathcal{F}})$ and $\operatorname{Aut}^{+}(\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}}),\psi_{\scriptscriptstyle\mathcal{G}})$ (see [10, Section 4]), the latter algebra is non-zero and hence $\operatorname{Aut}^{+}(\operatorname{C}(\mathbb{V}_{\mathcal{F}}),\psi_{\mathcal{F}})$ and $\operatorname{Aut}^{+}(\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}}),\psi_{\scriptscriptstyle\mathcal{G}})$ are monoidally equivalent. Then, the result follows from [10, Theorem 4.7]. ∎
@@
$p=[p_{ij}]_{i,j}\in\operatorname{B}\bigl{(}L^{2}(\mathcal{G}),L^{2}(\mathcal{F})\bigr{)}\otimes\mathcal{O}\bigl{(}\operatorname{QAut}(\mathcal{G}),\operatorname{QAut}(\mathcal{F})\bigr{)}$
be the unitary matrix defining $\mathcal{O}(\operatorname{QAut}(\mathcal{G}),\operatorname{QAut}(\mathcal{F}))$ as in [6, Definition 4.1] (remember the latter is non-zero) and let
$\rho_{\mathcal{F},\mathcal{G}}\colon\operatorname{C}(\mathcal{G})\ni e_{j}\longmapsto\sum\limits_{i}f_{i}\otimes p_{ij}\in\operatorname{C}(\mathcal{F})\otimes\mathcal{O}\bigl{(}\operatorname{QAut}(\mathcal{G}),\operatorname{QAut}(\mathcal{F})\bigr{)}$
be the corresponding unital *-homomorphism intertwining the quantum adjacency matrices. By Proposition 2.1 we know that $\delta_{\scriptscriptstyle\mathcal{G}}=\delta_{\mathcal{F}}$ and for the remainder of the proof we will denote this value by $\delta$. The maps introduced above can be represented as
$p\colon L^{2}(\mathcal{G})\otimes\mathcal{O}\bigl{(}\operatorname{QAut}(\mathcal{G}),\operatorname{QAut}(\mathcal{F})\bigr{)}\ni\xi\otimes a\longmapsto\sum\limits_{i,j}{\left|f_{i}\right\rangle}{\left\langle{e_{j}}\,\vline\,\xi\right\rangle}\otimes p_{ij}a\in L^{2}(\mathcal{F})\otimes\mathcal{O}\bigl{(}\operatorname{QAut}(\mathcal{G}),\operatorname{QAut}(\mathcal{F})\bigr{)}$

### v2

@@ Abstract.
We investigate a quantum generalization of the Mycielski construction for quantum graphs. In particular, we analyze the quantum symmetries of quantum Mycielskians and their relation to the symmetries of the underlying quantum graphs. We introduce a quantum analogue of graphs with twin vertices and show that, for classical graphs with a small number of vertices, this notion reduces to the classical one. Finally, we propose a quantum extension of the distinguishing number and examine its behavior under the quantum Mycielski construction.
@@ 1. Introduction
Classical graph theory has a wide range of intriguing applications beyond pure mathematics, with particularly strong connections to computer science and information theory. In communication theory, graphs are used to model channels, where vertices represent possible inputs and outputs, and edges encode confusability relation between signals. This perspective naturally extends to quantum information theory, where the role of confusability graphs is played by non-commutative operator systems associated with quantum channels (see, e.g., [7, 8, 13, 15, 19]). In this setting, Kraus operators of a quantum channel generate an algebraic structure that serves as the quantum analogue of the adjacency relations of a graph. The passage from classical to quantum thus requires reformulating standard graph invariants in operator-algebraic terms. Among the classical quantities that admit such generalizations are chromatic numbers - capturing the minimal resources needed for proper labelings - and clique numbers, which quantify the largest completely connected substructures. The (classical) chromatic number of a graph is the minimum number of labels (colors) that can be assigned to its vertices, so that the endpoints of every edge receive distinct labels. The clique number, by contrast, is the maximum size of a complete graph that can be embedded as a subgraph of the original. For example, graphs with clique number $2$ are precisely the triangle-free graphs. These invariants, when reinterpreted in the non-commutative framework, provide valuable insights into zero-error communication capacities of quantum channels and the interplay between combinatorial and operator-algebraic methods.
A long-standing problem in classical graph theory was to determine whether triangle-free graphs with arbitrarily large chromatic number exist. Equivalently, one asked whether the absence of small cliques imposes an upper bound on the chromatic complexity of a graph. This problem was resolved in [18], where Mycielski introduced an explicit iterative construction. Given a triangle-free graph $G$, he defined a new graph $\mu(G)$, now known as the Mycielskian of $G$, which is also triangle-free, contains $G$ as an induced subgraph, and whose chromatic number satisfies
$\chi(\mu(G))=\chi(G)+1.$
By iterating this construction, one obtains an infinite family of triangle-free graphs with unbounded chromatic number, thereby providing a negative answer to the aforementioned problem. The Mycielski construction has since become a paradigmatic example of how graph transformations can alter chromatic properties without affecting clique structure.
The natural question of extending this phenomenon to the quantum setting was addressed in [6], where two of the present authors introduced a non-commutative analogue of Mycielski’s construction. In this framework, classical graphs are replaced by quantum graphs, modeled by operator systems generated by Kraus operators of quantum channels. The proposed construction, termed the quantum Mycielskian, takes a finite quantum graph $\mathcal{G}$ as input and produces a new one that contains the original as a subgraph in the operator-algebraic sense. This quantum analogue preserves certain structural features of the classical case but exhibits important differences in the behavior of invariants. Specifically, it was shown that while the clique number remains unchanged under the transformation – mirroring the classical situation – the (quantum) chromatic number $\chi_{\bullet}$ does not necessarily increase by exactly one. Instead, it satisfies only a weaker property:
$\chi_{\bullet}(\mu(\mathcal{G}))\leq\chi_{\bullet}(\mathcal{G})+1,$
with strict equality not guaranteed in general.
In this paper, we investigate further properties of the Mycielski construction in the quantum framework, with an emphasis on how structural and symmetry-related features behave under this transformation. Section 1.1 begins with a review of the classical Mycielski construction, highlighting its role in generating triangle-free graphs of arbitrarily large chromatic number. Section 1.2 then sets up the operator-algebraic formalism of quantum graphs and introduces the actions of quantum groups on them. These actions provide the natural symmetry framework in the non-commutative setting, and their behavior under graph transformations will play a central role in our analysis.
A key technical aspect of this study concerns actions that preserve partitions of unity, which, in the commutative case, correspond precisely to proper vertex colorings or labelings of graphs. We discuss this notion in Section 1.3, and its relevance for quantum analogues of chromatic invariants will be revisited throughout the paper (see, in particular, Section 4). At the end of Section 1, we recall the definition of the quantum Mycielskian and prepare the ground for our new results.
Section 2 is devoted to isomorphisms of quantum Mycielskians. In Section 3, we investigate how the Mycielski construction interacts with the quantum automorphism groups of graphs, showing that this process naturally gives rise to the notion of quantum twin vertices. These are studied in detail in Sections 3.1 and 3.2, as well as in the Appendix.
Finally, in Section 4, we introduce the quantum distinguishing number, a non-commutative analogue of the classical distinguishing number that measures the minimal labeling required to break all symmetries of a graph. We analyze how this invariant behaves under the quantum Mycielski construction and compare the results with the classical situation, further illustrating both the parallels and the divergences between the two frameworks.
@@ 1.1. Classical Mycielski construction
Let $G=(V_{G},E_{G})$ be a (finite, undirected) graph. The Mycielskian of $G$, denoted by $\mu(G)$ is a graph with the set of vertices $V_{\mu(G)}=\{\bullet\}\sqcup V_{G}\sqcup V_{G}$, where $\bullet$ denotes a distinguished point, called master vertex. The edge set $E_{\mu(G)}$ can be specified by the adjacency matrix $A_{\scriptscriptstyle\mu(G)}$. We recall that the adjacency matrix $A_{\scriptscriptstyle G}$ for an undirected graph $G$ is a $\{0,1\}$-matrix whose columns and rows are indexed by the set of vertices, and $(A_{\scriptscriptstyle G})_{i,j}$ is non-zero if and only if $\{v_{i},v_{j}\}\in E$. For the Mycielskian, we have
$A_{\scriptscriptstyle\mu(G)}=\begin{bmatrix}0&\vec{0}^{T}&\vec{1}^{T}\\ \vec{0}&A_{\scriptscriptstyle G}&A_{\scriptscriptstyle G}\\ \vec{1}&A_{\scriptscriptstyle G}&0\end{bmatrix}$,
where $\vec{1}$ (resp. $\vec{0}$) is the column vector with all entries being $1$ (resp. 0). Mycielski’s theorem [18] says that denoting by $\omega(\cdot)$ the clique number of a graph we have $\omega(\mu(G))=\omega(G)$, i.e. the clique number remains unchanged under the passage from $G$ to $\mu(G)$, and for the chromatic numbers $\chi$ we have $\chi(\mu(G))=\chi(G)+1$.
@@
$V_{\mu_{r-1}(G)}=\{\bullet\}\sqcup V^{1}\sqcup\ldots\sqcup V^{r}$
@@
$E=E_{G}\cup\bigcup_{i=0}^{r-1}\bigl\{\{v^{i}_{j},v^{i+1}_{j^{\prime}}\}\,\bigr|\bigl.\,\{v_{j},v_{j^{\prime}}\}\in E_{G}\bigr\}\cup\bigl\{\{v^{r}_{1},\bullet\},\{v^{r}_{2},\bullet\},\ldots,\{v^{r}_{|V_{G}|},\bullet\}\bigr\}.$
@@ 1.2. Quantum graphs and their quantum symmetries
Let $\mathfrak{C}$ be the category of $\mathrm{C}^{*}$-algebras as defined in [27, 25], and $\mathfrak{C}^{\mathrm{op}}$ its dual category. A quantum space $\mathbb{X}$ is an object from $\mathfrak{C}^{\mathrm{op}}$ and its corresponding $\mathrm{C}^{*}$-algebra is denoted by $\operatorname{C}(\mathbb{X})$. We say that a quantum space $\mathbb{X}$ is finite if the $\mathrm{C}^{*}$-algebra $\operatorname{C}(\mathbb{X})$ is finite-dimensional.
@@
For alternative approaches to quantum graphs and comparison between them see [11] and references therein. In what follows, the $\mathrm{C}^{*}$-algebra $\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}})$ will be denoted by $\operatorname{C}(\mathcal{G})$ and the Hilbert space $L^{2}(\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}}),\psi_{\scriptscriptstyle\mathcal{G}})$ by $L^{2}(\mathcal{G})$. We will also shorten $\delta(\psi_{\scriptscriptstyle\mathcal{G}})$ to $\delta_{\scriptscriptstyle\mathcal{G}}$. A quantum graph $\mathcal{G}$ is called irreflexive if $m(A\otimes\mathrm{id}_{\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}})})m^{\ast}=0$.
Definition 1.2 ([8, Definition 6.3]).
@@
The quantum symmetry group of a quantum graph $\mathcal{G}$ is by construction a quantum subgroup of a well-studied quantum group $\operatorname{Aut}^{+}(\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}}),\psi_{\scriptscriptstyle\mathcal{G}})$ of symmetries of the quantum space $\mathbb{V}_{\scriptscriptstyle\mathcal{G}}$ together with the $\delta$-form $\psi_{\scriptscriptstyle\mathcal{G}}$ (the action of this larger quantum group does not preserve the adjacency matrix). The quantum groups of symmetries of finite quantum spaces equipped with a $\delta$-form were introduced in [3]. Finally, we note that in the existing literature the notion of an action of quantum groups is often referred to as a coaction of the corresponding $\mathrm{C}^{*}$-algebras. For brevity, however, we adopt the former convention.
@@ 1.3. Actions preserving a partition of unity
Let $\mathsf{A}$ be a unital $\mathrm{C}^{*}$-algebra and let $\rho\colon\mathsf{A}\to\mathsf{A}\otimes\operatorname{C}(\mathbb{G})$ be an action of a compact quantum group $\mathbb{G}$ on the quantum space underlying $\mathsf{A}$. In the subsequent sections of this paper, $\mathsf{A}$ will often take the form $\operatorname{C}(\mathbb{X})$ or $\operatorname{C}(\mathbb{X})\otimes\operatorname{B}(\mathscr{H})$, for some Hilbert space $\mathscr{H}$, depending on the context. Let $\mathfrak{P}=\{P_{1},\dotsc,P_{n}\}$ be a partition of unity in $\mathsf{A}$. Let $\mathsf{J}$ be the ideal in $\operatorname{C}(\mathbb{G})$ generated by the elements
@@
$\begin{aligned} (\Delta_{\mathbb{G}_{\scriptscriptstyle{\mathfrak{P}}}}\otimes\mathrm{id})\circ\Delta_{\mathbb{G}_{\scriptscriptstyle{\mathfrak{P}}}}\circ\pi&=(\Delta_{\mathbb{G}_{\scriptscriptstyle{\mathfrak{P}}}}\otimes\mathrm{id})\circ(\pi\otimes\pi)\circ\Delta_{\mathbb{G}}=(\pi\otimes\pi\otimes\pi)\circ(\Delta_{\mathbb{G}}\otimes\mathrm{id})\circ\Delta_{\mathbb{G}}\\ &=(\pi\otimes\pi\otimes\pi)\circ(\mathrm{id}\otimes\Delta_{\mathbb{G}})\circ\Delta_{\mathbb{G}}=(\mathrm{id}\otimes\Delta_{\mathbb{G}_{\scriptscriptstyle{\mathfrak{P}}}})\circ(\pi\otimes\pi)\circ\Delta_{\mathbb{G}}=(\mathrm{id}\otimes\Delta_{\mathbb{G}_{\scriptscriptstyle{\mathfrak{P}}}})\circ\Delta_{\mathbb{G}_{\scriptscriptstyle{\mathfrak{P}}}}\circ\pi,\end{aligned}$
@@ 1.4. Mycielskians of quantum graphs
Here we recall the construction introduced in [6]. For $r\geq{2}$ the Mycielski construction of order r1r-1 for quantum graphs is the procedure of passing from a quantum graph $\mathcal{G}$ to a new quantum graph $\mu_{r-1}(\mathcal{G})$ defined as follows ([6, Section 4]):
@@
    $A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\left[\begin{smallmatrix}\lambda\\ x_{1}\\[-2.0pt] \vdots\\[2.0pt] x_{r}\end{smallmatrix}\right]=\left[\begin{smallmatrix}\delta_{\scriptscriptstyle\mathcal{G}}^{2}\psi_{\scriptscriptstyle\mathcal{G}}(x^{r})\\ A_{\scriptscriptstyle\mathcal{G}}(x_{1}+x_{2})\\ A_{\scriptscriptstyle\mathcal{G}}(x_{1}+x_{3})\\[-2.0pt] \vdots\\[2.0pt] A_{\scriptscriptstyle\mathcal{G}}(x_{r-2}+x_{r})\\ \lambda\mathds{1}_{\operatorname{C}(\mathcal{G})}+A_{\scriptscriptstyle\mathcal{G}}x_{r-1}\end{smallmatrix}\right].$
@@
for $k=1,\dotsc,r$. Note that although $\iota_{0},\dotsc,\iota_{r}$, and their adjoints are not multiplicative, they are all *-maps. In this notation, we have (cf. [6, Eq. (9)]):
$A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}=\delta_{\scriptscriptstyle\mathcal{G}}\iota_{r}\eta_{\scriptscriptstyle\mathcal{G}}\iota_{0}^{\ast}+\delta_{\scriptscriptstyle\mathcal{G}}\iota_{0}\eta_{\scriptscriptstyle\mathcal{G}}^{\ast}\iota_{r}^{\ast}+\iota_{1}A_{\scriptscriptstyle\mathcal{G}}\iota_{1}^{\ast}+\sum\limits_{k=1}^{r-1}\bigl(\iota_{k}A_{\scriptscriptstyle\mathcal{G}}\iota_{k+1}^{\ast}+\iota_{k+1}A_{\scriptscriptstyle\mathcal{G}}\iota_{k}^{\ast}\bigr).$
@@ 2. Quantum isomorphisms of Mycielskians
Let $\mathcal{G}$ and $\mathcal{F}$ be quantum graphs. The notion of quantum isomorphism of $\mathcal{G}$ and $\mathcal{F}$ has been studied in [7, Definition 4.4, Corollary 4.8] and [8, Definition 6.5, Theorem 6.6]. The definitions of quantum isomorphism of quantum graphs given in the two papers are equivalent ([8, Theorem 6.6]) and we will use both formulations.
According to [8, Definition 6.5] we say that $\mathcal{G}$ and $\mathcal{F}$ are quantum isomorphic if there is a monoidal equivalence of the quantum automorphism groups $\operatorname{Aut}^{+}(\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}}),\psi_{\scriptscriptstyle\mathcal{G}})$ and $\operatorname{Aut}^{+}(\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{F}}),\psi_{\scriptscriptstyle\mathcal{F}})$ which maps the monoid object $\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}})$ to the monoid object $\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{F}})$ and maps the morphism $A_{\scriptscriptstyle\mathcal{G}}$ to $A_{\scriptscriptstyle\mathcal{F}}$ (the notion of monoidal equivalence of compact quantum groups is described e.g. in [12]).
The equivalent formulation of quantum isomorphism of $\mathcal{G}$ and $\mathcal{H}$ given in [7, Section 4] involves the *-algebra $\mathcal{O}(\operatorname{QAut}(\mathcal{F}),\operatorname{QAut}(\mathcal{G}))$ generated by elements $\{p_{ij}\}$ of a unitary matrix with the relations ensuring that upon a choice of orthonormal bases $\{e_{j}\}$ and $\{f_{i}\}$ of $\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}})$ and $\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{F}})$ the map
$\rho_{\scriptscriptstyle\mathcal{F},\mathcal{G}}\colon e_{i}\longmapsto\sum_{i}f_{i}\otimes p_{ij}$
is a unital *-homomorphism $\operatorname{C}(\mathcal{G})\to\operatorname{C}(\mathcal{F})\otimes\mathcal{O}(\operatorname{QAut}(\mathcal{F}),\operatorname{QAut}(\mathcal{G}))$ such that $\rho_{\scriptscriptstyle\mathcal{F},\mathcal{G}}A_{\scriptscriptstyle\mathcal{G}}=(A_{\scriptscriptstyle\mathcal{F}}\otimes\mathrm{id})\rho_{\scriptscriptstyle\mathcal{F},\mathcal{G}}$. The quantum graphs $\mathcal{G}$ and $\mathcal{F}$ are isomorphic if and only if the algebra $\mathcal{O}(\operatorname{QAut}(\mathcal{F}),\operatorname{QAut}(\mathcal{G}))$ is non-zero. Note that the matrix $p=[p_{ij}]_{i,j}$ can be treated as an element of $\operatorname{B}(L^{2}(\mathcal{G}),L^{2}(\mathcal{F}))\otimes\mathcal{O}(\operatorname{QAut}(\mathcal{F}),\operatorname{QAut}(\mathcal{G}))$.
@@
Let $\mathcal{F}$ and $\mathcal{G}$ be quantum graphs. If $\mathcal{F}$ and $\mathcal{G}$ are quantum isomorphic, then $\delta_{\scriptscriptstyle\mathcal{F}}=\delta_{\scriptscriptstyle\mathcal{G}}$.
@@
According to the first version of the definition of quantum isomorphism of $\mathcal{G}$ and $\mathcal{F}$ given above, the quantum groups $\operatorname{Aut}^{+}(\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{F}}),\psi_{\scriptscriptstyle\mathcal{F}})$ and $\operatorname{Aut}^{+}(\operatorname{C}(\mathbb{V}_{\scriptscriptstyle\mathcal{G}}),\psi_{\scriptscriptstyle\mathcal{G}})$ are monoidally equivalent. Hence the result follows from [12, Theorem 4.7]. ∎
@@
$p=[p_{ij}]_{i,j}\in\operatorname{B}\bigl(L^{2}(\mathcal{G}),L^{2}(\mathcal{F})\bigr)\otimes\mathcal{O}\bigl(\operatorname{QAut}(\mathcal{F}),\operatorname{QAut}(\mathcal{G})\bigr)$
be the unitary matrix defining $\mathcal{O}(\operatorname{QAut}(\mathcal{F}),\operatorname{QAut}(\mathcal{G}))$ as above (remember the latter is non-zero) and let
$\rho_{\scriptscriptstyle\mathcal{F},\mathcal{G}}\colon\operatorname{C}(\mathcal{G})\ni e_{j}\longmapsto\sum\limits_{i}f_{i}\otimes p_{ij}\in\operatorname{C}(\mathcal{F})\otimes\mathcal{O}\bigl(\operatorname{QAut}(\mathcal{F}),\operatorname{QAut}(\mathcal{G})\bigr)$
be the corresponding unital *-homomorphism intertwining the quantum adjacency matrices. By Proposition 2.1 we know that $\delta_{\scriptscriptstyle\mathcal{G}}=\delta_{\scriptscriptstyle\mathcal{F}}$ and for the remainder of the proof we will denote this value by $\delta$. The maps introduced above can be represented as
$p\colon L^{2}(\mathcal{G})\otimes\mathcal{O}\bigl(\operatorname{QAut}(\mathcal{F}),\operatorname{QAut}(\mathcal{G})\bigr)\ni\xi\otimes a\longmapsto\sum\limits_{i,j}{\left|f_{i}\right\rangle}{\left\langle{e_{j}}\,\vline\,\xi\right\rangle}\otimes p_{ij}a\in L^{2}(\mathcal{F})\otimes\mathcal{O}\bigl(\operatorname{QAut}(\mathcal{F}),\operatorname{QAut}(\mathcal{G})\bigr)$